OpenStudy (anonymous):

Write the standard equation for the circle. center (10, –6), r = 6 A. (x + 10)2 + (y – 6)2 = 6 B. (x + 6)2 + (y – 10)2 = 36 C. (x – 10)2 + (y + 6)2 = 36 D. (x – 10)2 + (y + 6)2 = 6

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2 ( this is the standard equation of circle ) where a , b is the point on the circle and r is radius... hope u get your answer now

OpenStudy (anonymous):

C. (x – 10)2 + (y + 6)2 = 36